### v2

# FCS-HGNN: Flexible Multi-type Community Search in Heterogeneous Information Networks

Abstract.
Community search is a personalized community discovery problem designed to identify densely connected subgraphs containing the query node. Recently, community search in heterogeneous information networks (HINs) has received considerable attention. Existing methods typically focus on modeling relationships in HINs through predefined meta-paths or user-specified relational constraints. However, metapath-based methods are primarily designed to identify single-type communities with nodes of the same type rather than multi-type communities involving nodes of different types. Constraint-based methods require users to have a good understanding of community patterns to define a suitable set of relational constraints, which increases the burden on users. In this paper, we propose FCS-HGNN, a novel method for flexibly identifying both single-type and multi-type communities in HINs. Specifically, FCS-HGNN extracts complementary information from different views and dynamically considers the contribution of each relation instead of treating them equally, thereby capturing more fine-grained heterogeneous information. Furthermore, to improve efficiency on large-scale graphs, we further propose LS-FCS-HGNN, which incorporates i) the neighbor sampling strategy to improve training efficiency, and ii) the depth-based heuristic search strategy to improve query efficiency. We conducted extensive experiments to demonstrate the superiority of our proposed methods over state-of-the-art methods, achieving average improvements of 14.3% and 11.1% on single-type and multi-type communities, respectively.
1. Introduction
Community search, a kind of query-dependent personalized community discovery problem designed to identify densely-connected subgraphs containing query nodes, has garnered considerable interest (Luo et al., 2020; Guo et al., 2021; Chen et al., 2023b; Cai et al., 2023; Chen and Gao, 2022; Miao et al., 2022; Wang et al., 2024; Chen et al., 2023a). Most existing studies primarily focus on homogeneous graphs associated with a single type of nodes and edges. However, many real-world data are naturally represented as heterogeneous information networks (HINs), where multiple types of entities and the relations between them are embodied by various types of nodes and edges, respectively. Compared with homogeneous graphs, HINs possess richer semantics and complex relationships, encapsulating more valuable community information.
Recent advances in community search over HINs usually endeavor to model the relationships between nodes of different types through the predefined meta-paths (Fang et al., 2020; Qiao et al., 2021; Jiang et al., 2022a; Zhou et al., 2023; Liu et al., 2023) or user-specified relational constraints (Jian et al., 2020). However, as shown in Table 1, they suffer from four main limitations. (1) These metapath-based methods are designed to identify single-type communities with nodes of the same type, rather than multi-type communities involving nodes of different types. The single-type community does not fully exploit the valuable community information in HINs. For example, as shown in Figure 1, the single-type community composed of authors (blue circle) can only uncover collaboration information among authors. Instead, a multi-type community spanning authors, papers, and venues (red circle) may reflect the core of a research field. By analyzing such multi-type communities, we can not only understand collaboration patterns among authors but also delve deeper into research trends by papers and influential venues in this field, thereby more effectively supporting various downstream applications (Chakraborty et al., 2015; Li et al., 2018b; Gajewar and Das Sarma, 2012; Behrouz et al., 2022). (2) Constraint-based methods require users to have a good understanding of community patterns, in order to define a suitable set of relational constraints. Otherwise, without any guidance, they may fail to find any communities, which significantly increases the burden on end-users. (3) Both metapath-based and constraint-based methods suffer from the pattern inflexibility, which refers to the problem that they are based on predefined community patterns, such as $(k,\mathcal{P})$-core (Fang et al., 2020; Zhou et al., 2023; Liu et al., 2023), and $(k,\Psi)$-NMC (Jiang et al., 2022a), or the user-specified relational constraints. In real-world scenarios, the pattern of a community is flexible in nature, and the target communities may not conform to predefined community patterns or relational constraints, thus risking being overlooked. (4) Both metapath-based and constraint-based methods only focus on structural properties of HINs (Fang et al., 2020; Jian et al., 2020; Jiang et al., 2022a) or limited low-dimensional node features (Zhou et al., 2023; Liu et al., 2023). For example, (Zhou et al., 2023) only considers one-dimensional features (such as influence). In practical scenarios, HINs usually naturally endow nodes with high-dimensional features containing important information, which is a challenge for existing methods in HINs.
Table 1. Comparison with other CS Methods in HINs.
Method
Single
type
Multi
type
Node
feature
Meta-path
Community
patterns
e.g., (Fang et al., 2020; Jiang et al., 2022a) Predefined Predefined
e.g., (Zhou et al., 2023; Liu et al., 2023) ✓(Low Dim.) Predefined Predefined
e.g., (Jian et al., 2020) Not Dependent User-specified
Ours Not Dependent Data driven
Figure 1. The blue and red circles represent single-type and multi-type community examples in the bibliographic network, respectively.
To address the above limitations, we propose FCS-HGNN, a query-driven heterogeneous graph neural network that can flexibly identify multi-type communities, as well as single-type communities in HINs. To the best of our knowledge, we are the first to propose a machine learning (ML) based community search algorithm tailored for heterogeneous information networks. To tackle pattern inflexibility, FCS-HGNN adaptively learns community patterns in a data-driven manner rather than relying on strict predefined community patterns or user-specified relational constraints. Specifically, we design two encoders: (1) Heterogeneous Encoder to deal with the heterogeneity and learn query-independent node features, and (2) Query Encoder to learn the local structural information around the query nodes. In each encoder, we introduce an edge semantic attention mechanism to dynamically consider the contribution of each relation when identifying diverse multi-type communities rather than treating them equally. It is worth noting that graph neural networks offer a unified way to combine the structural information and the arbitrary dimensional node features. To take advantage of complementary information from different encoders, FCS-HGNN integrates them through an attention mechanism after each layer and ultimately utilizes them in the final layer to infer the probability of each node belonging to the community. Finally, by considering the nodes’ probabilities belonging to the community, we employ a breadth-first search (BFS) based algorithm to obtain the target multi-type community.
Furthermore, to improve efficiency on large-scale graphs, we introduce the neighbor sampling algorithm and depth-based heuristic search strategy to further optimize our algorithm, named LS-FCS-HGNN (Large Scale FCS-HGNN). Through the neighbor sampling algorithm, in each iteration, LS-FCS-HGNN updates its parameters on a randomly sampling subgraph, rather than on the entire large-scale graph, which significantly reduces the computational resource requirements and improves the efficiency during the training phase. Furthermore, based on real-life observations, we find that the nodes within the same community typically exhibit more edges and shorter paths. In other words, the nodes closer to the query node are more likely to belong to the community. Therefore, we propose a depth-based heuristic search strategy to reduce searches for the nodes farther from the query node, thereby significantly improving query efficiency.
To summarize, we make the following contributions.
$\bullet$ We propose FCS-HGNN, a novel query-driven heterogeneous graph neural network capable of effectively identifying multi-type communities, as well as single-type communities in heterogeneous information networks. FCS-HGNN adaptively learns community patterns in a data-driven manner, without relying on predefined community patterns and user-specified relational constraints, significantly alleviating the burden on end-users.
$\bullet$ We further propose LS-FCS-HGNN, an advanced algorithm designed to improve the efficiency in large-scale graphs. LS-FCS-HGNN incorporates two strategies including i) the neighbor sampling algorithm to reduce the computational resource requirements and improve training efficiency, and ii) the depth-based heuristic strategy to improve the query efficiency, while preserving community effectiveness.
$\bullet$ We conduct extensive experiments on five real-world heterogeneous information networks to demonstrate the superiority of our proposed methods over state-of-the-art methods. Our proposed FCS-HGNN and LS-FCS-HGNN significantly enhance the effectiveness of multi-type community search and improve the training and query efficiency.
Roadmap. The rest of the paper is organized as follows. Section 2 discusses related work. Section 3 provides some preliminaries. Section 4 details our proposed FCS-HGNN. Section 5 elaborates on our advanced LS-FCS-HGNN. Section 6 conducts comprehensive experiments to demonstrate the effectiveness of our proposed methods. Finally, section 7 concludes the paper.
2. RELATED WORK
Community Search. Over the past few decades, there has been extensive research on community search, with the primary objective of identifying a densely connected subgraph that encompasses the query nodes. To measure the structural cohesiveness of communities, various community patterns have been proposed, such as $k$-core (Seidman, 1983; Batagelj and Zaversnik, 2003), $k$-truss (Cohen, 2008; Huang et al., 2014), $k$-clique (Tsourakakis, 2015), and $k$-edge connected component (ECC) (Chang et al., 2015; Hu et al., 2016). However, the patterns of communities in real-world scenarios are flexible in nature. If real-world communities do not follow any of the above patterns, these models will fail to discover the true communities. Therefore, machine learning (ML)-based community search methods have been proposed (Gao et al., 2021; Jiang et al., 2022b; Li et al., 2023; Fang et al., 2023; Behrouz and Hashemi, 2022; Hashemi et al., 2023) to address the limitation of pattern inflexibility. However, most of these works primarily focus on homogeneous graphs where nodes belong to the same type.
Recently, an increasing number of researchers have shown interest in community search over heterogeneous information networks (HINs) (Fang et al., 2021, 2020; Jian et al., 2020; Qiao et al., 2021; Jiang et al., 2022a; Zhou et al., 2023; Liu et al., 2023), due to the prevalence and significance of HINs in real-world scenarios. These studies usually model the relationships between nodes in HINs through predefined meta-paths or user-specified relational constraints. For example, (Fang et al., 2020) proposes to employ predefined meta-paths to delineate the connectivity among nodes of different types in HINs, and extend the $k$-core to $(k,\mathcal{P})$-core to measure the cohesiveness of a community, where $\mathcal{P}$ is the predefined meta-path. Metapath-based methods aim to find single-type communities consisting of nodes with the same type, rather than discovering multi-type communities involving nodes of different types. (Jian et al., 2020) proposes a relational community model, which defines communities based on a set of user-specified relational constraints. It requires users to have a good understanding of community patterns to define a suitable set of constraints, which significantly increases the burden on users. If the constraints are not specified correctly, it may fail to find any communities. In contrast, our FCS-HGNN alleviates pattern inflexibility by adaptively learning community patterns in a data-driven manner, thereby alleviating the burden on end-users.
Heterogeneous Graph Representation Learning. A large amount of work (Wang et al., 2022) has been devoted to address the problem of learning node representations from HINs, including heterogeneous graph embedding (Dong et al., 2017; Fu et al., 2017) and heterogeneous graph neural networks (Wang et al., 2019; Zhang et al., 2019; Lv et al., 2021; Yang et al., 2023). Heterogeneous graph embedding primarily focuses on preserving structure information based on meta-paths. For example, metapath2vec (Dong et al., 2017) devised a metapath-based random walk and utilized skip-gram (Mikolov et al., 2013) to learn node representations. Furthermore, according to the way of dealing with heterogeneity, heterogeneous graph neural networks can be divided into metapath-based methods and metapath-free methods. The design and selection of meta-paths still pose challenges in practice, which may require sufficient domain knowledge. When identifying multi-type communities, predefined meta-paths may lack flexibility, as different meta-paths are required for communities involving different user-specified node types. In contrast, our proposed FCS-HGNN adopts a metapath-free manner (Heterogeneous Encoder) to overcome heterogeneity and employs the Query Encoder to learn local structural information around query nodes, which sufficiently facilitates the learning of node representations in HINs.
3. Preliminaries
In this section, we first define heterogeneous information networks (HINs) and the multi-type community. Then, we formalize the multi-type community search problem in HINs.
Definition 0 (Heterogeneous Information Network).
A heterogeneous information network is a directed graph $\mathcal{H}=(V,E,\psi,\phi)$ with a node type mapping function $\psi:V\rightarrow\mathcal{A}$ and an edge type mapping function $\phi:E\rightarrow\mathcal{R}$, where $V$ and $E$ represent the sets of nodes and edges respectively, each node $v\in V$ belongs to a node type $\psi(v)\in\mathcal{A}$, each edge $e\in E$ belongs to an edge type $\phi(e)\in\mathcal{R}$, and $|\mathcal{A}|+|\mathcal{R}|>2$.
Due to its ability to delineate multiple types of nodes and edges as well as complex relationships, HINs have become ubiquitous in real-world scenarios, such as bibliographic networks (Sun et al., 2011; Shi et al., 2012a), IMDB movie networks (Shi et al., 2012b; Yu et al., 2013), and knowledge graphs (Hogan et al., 2021). Compared to homogeneous graphs, HINs encapsulate richer community information. Most existing methods primarily focus on single-type communities by meta-paths, which may not fully explore valuable community information in HINs. Therefore, we propose the multi-type community involving nodes of different types as follows:
Definition 0 (Multi-type Community).
Given a HIN $\mathcal{H}=(V,E,\psi,\phi)$ and the set of node types $\mathcal{S}_{\mathcal{A}}\subseteq\mathcal{A}$, a multi-type community is defined as a subgraph $\mathcal{C}_{q}=(V_{q},E_{q},\psi,\phi)$ of $\mathcal{H}$ that satisfies the following criteria:
  1. (1)
    Each node $v\in V_{q}$ satisfies $\psi(v)\in\mathcal{S}_{\mathcal{A}}$;
  2. (2)
    $\mathcal{C}_{q}$ is connected and has a cohesive structure.
According to definition 2, the single-type community is a special case (when $|\mathcal{S}_{\mathcal{A}}|=1$) of the multi-type community. Metapath-based methods can identify single-type communities by predefined meta-paths, while constraint-based methods can identify multi-type communities by defining a set of relational constraints, as shown in Table 1. In contrast, our proposed FCS-HGNN can flexibly identify both single-type communities and multi-type communities without relying on strictly predefined meta-paths or relational constraints. Therefore, we formalize the problem statement for multi-type community search in HINs.
Problem 1 (Multi-type Community Search in HINs).
Given a HIN $\mathcal{H}=(V,E,\psi,\phi)$, a query node $v_{q}\in V$, and the set of node types $\mathcal{S}_{\mathcal{A}}$, multi-type community search aims to identify the query-dependent community $\mathcal{C}_{q}=(V_{q},E_{q},\psi,\phi)$ satisfying Multi-type Community (refer to Definition 2) and $v_{q}\in V_{q}$.
Based on the above problem definition, we formalized the machine learning-based training manner as follows: given the set of node types $\mathcal{S}_{\mathcal{A}}$ and a set of query tasks with labeled communities, denoted as $\mathcal{D}=\{(q_{1},pos(q_{1}),neg(q_{1})),(q_{2},pos(q_{2}),neg(q_{2})),\cdots\}$, where $q$ represents a query node, $pos(q)$ is the set of positive nodes in the target community, $neg(q)$ is the set of negative nodes not in the target community. Our training objective is to learn a function $\mathcal{F}(\mathcal{H},q,\mathcal{S}_{\mathcal{A}})$ that yields the probability of each node belonging to the target community. Then, we can employ BFS to find the target community with the highest probability. It is worth noting that $neg(q)$ can be randomly sampled from the nodes except for the query node and $pos(q)$, without incurring additional annotation costs.
4. FCS-HGNN
In this section, we first provide an overview of the FCS-HGNN, followed by a detailed description of each component in FCS-HGNN.
4.1. Framework
Figure 2. The overall framework of FCS-HGNN.
We illustrate the overall framework of FCS-HGNN in Figure 2, which includes three main steps: Heterogeneous and Query Information Preprocessing, Information Integration, and Search Algorithm. In heterogeneous and query information preprocessing, since different types of nodes in HINs usually have different semantic features, we design a type-shared feature projection to map different types of features to a unified feature space, thereby obtaining a new graph denoted as $\mathcal{H}_{h}$. In addition, we incorporate query-dependent information into HIN, denoted $\mathcal{H}_{q}$, where the graph structure maintains consistency with $\mathcal{H}$. To better utilize information from different views provided by $\mathcal{H}_{h}$ and $\mathcal{H}_{q}$, we design two encoders in information integration: the heterogeneous encoder (abbreviated as ”Hetero. Encoder” in Figure 2) and the query encoder. In each encoder, we propose an edge semantic attention mechanism to selectively aggregate neighbor information based on different types of relations, thereby capturing more fine-grained heterogeneous information. Meanwhile, FCS-HGNN integrates complementary information (query-dependent and query-independent information) after each layer using an attention mechanism. Ultimately, FCS-HGNN utilizes this integrated information to predict the probability of each node belonging to the community. In the search algorithm, we search the target community by BFS based on the probability of each node and user-specified node types $\mathcal{S}_{\mathcal{A}}$.
4.2. Heterogeneous & Query Information Preprocessing
In the community search task over HINs, we primarily have to deal with two aspects of information: the query-dependent information provided by the user and the query-independent information inherent in HINs, such as the node features and the graph structure. Differing from homogeneous graphs, HINs typically comprise multiple types of nodes and edges, presenting significant challenges to community search due to their heterogeneity. Existing studies on community search usually deal with the heterogeneity among nodes through predefined meta-paths (Fang et al., 2020; Qiao et al., 2021; Jiang et al., 2022a; Zhou et al., 2023; Liu et al., 2023) or user-specified relationship constraints (Jian et al., 2020). However, metapath-based methods are designed to identify single-type communities with nodes of the same type rather than multi-type communities. On the other hand, constraint-based methods require users to have a thorough understanding of community patterns; otherwise, they may fail to identify any community.
4.2.1. Type-shared Feature Projection
To address these issues, we propose a type-shared feature projection to elegantly deal with the heterogeneity. In HINs, nodes of different types typically reside in distinct feature spaces, which hinders the information aggregation among them (Wang et al., 2019; Yang et al., 2023). Therefore, we assume the existence of a unified feature space $\mathcal{U}\in\mathbb{R}^{d_{u}}$. For each node type $t\in\mathcal{T}$, where $\mathcal{T}=\{\psi(v_{i})|v_{i}\in V\}$, we project the original feature of each node $x[v_{i}]$ to the unified feature space $\mathcal{U}$ through a learnable type-shared linear layer with the following formula:
(1) $\hat{x}[v_{i}]=W_{t}x[v_{i}]$
where $\forall v_{i}\in V$, $t$ represents the type of node $v_{i}$, i.e., $\psi(v_{i})=t$, $W_{t}\in\mathbb{R}^{d_{t}\times d_{u}}$ is a learnable and type-specific linear layer, $d_{t}$ denotes the feature dimension of the original space for node type $t$, $d_{u}$ denotes the feature dimension of the unified space $\mathcal{U}$.
Equation 1 enables nodes of different types to effectively aggregate information in a unified space, alleviating the burden on systems or users as they no longer need to design predefined meta-paths or a series of relationship constraints to handle heterogeneity. It is worth noting that we still preserve the heterogeneity of edges in $\mathcal{H}_{h}$ rather than disregarding it, and propose an edge semantic attention mechanism (elaborated in Section 4.3) to capture more fine-grained heterogeneous information.
4.2.2. Query-dependent Graph
In addition to the node feature and graph structure inherent in HINs, query information is also very crucial for community search. Therefore, we encode query-dependent information $\mathcal{X}_{q}$ into the HIN $\mathcal{H}$, denoted as $\mathcal{H}_{q}$, where the graph structure maintains consistency with $\mathcal{H}$. Here, $\mathcal{X}_{q}\in\mathbb{R}^{n\times 1}$ is associated with the query node $q$ and is composed of $x_{q}[v_{i}]$, $\forall{v_{i}}\in V$.
(2) $x_{q}[v_{i}]=\begin{dcases}1&\text{if }v_{i}=q;\\ 0&\text{if }v_{i}\neq q.\end{dcases}$
The query information $\mathcal{X}_{q}$ shares the same semantics, indicating whether they are query nodes, allowing nodes of different types to reside in the same feature space. However, the same as $\mathcal{H}_{h}$, we maintain the edge heterogeneity in $\mathcal{H}_{q}$, preserving the type of each edge. This is done to effectively preserve heterogeneous structural information around the query node during the information integration, rather than simply treating $\mathcal{H}_{q}$ as a homogeneous graph.
As shown in Figure 2, we obtain $\mathcal{H}_{h}$ and $\mathcal{H}_{q}$ from $\mathcal{H}$ by employing type-shared feature projection and introducing query-dependent information. $\mathcal{H}_{h}$ and $\mathcal{H}_{q}$ offer two distinct views containing complementary information. The former describes the features of nodes with various types in a unified space, which is beneficial to information extraction and integration in the subsequent steps. The latter delineates the local structural information around the query node, which is crucial for community search.
4.3. Information Integration
To better extract and integrate information from different views ($\mathcal{H}_{h}$ and $\mathcal{H}_{q}$), we designed two encoders: the heterogeneous encoder and the query encoder. First, we propose the edge semantic attention mechanism to aggregate neighbor information, which is the basis of these two encoders.
4.3.1. Edge Semantic Attention Mechanism
Different from homogeneous graphs, edges in HINs contain complex relationships, which are crucial for understanding the relationships between nodes and mining valuable community information. For example, in bibliography networks (Shi et al., 2012a; Sun et al., 2011), author-paper (A-P) edges represent authoring relationships, and paper-venues (P-V) edges represent publishing relationships. Therefore, the heterogeneity inherent in edges still requires careful consideration. Existing methods (Fang et al., 2020; Jian et al., 2020; Jiang et al., 2022a; Zhou et al., 2023; Liu et al., 2023) treat different types of relations equally when identifying communities. However, in multi-type communities, the contributions of different relations usually vary across different communities.
To address the above issues, we propose the edge semantic attention mechanism as follows:
(3) $\alpha_{ij}=\frac{\exp{\Big{(}\mathrm{LeakyReLU}\big{(}\vec{\mathbf{a}}^{T}[\mathbf{W}\vec{\mathbf{h}}_{i}\|\mathbf{W}\vec{\mathbf{h}}_{j}\|\mathbf{W_{e}}\vec{\mathbf{e}}_{ij}]\big{)}\Big{)}}}{\sum_{r\in\mathcal{N}_{i}}\exp{\Big{(}\mathrm{LeakyReLU}\big{(}\vec{\mathbf{a}}^{T}[\mathbf{W}\vec{\mathbf{h}}_{i}\|\mathbf{W}\vec{\mathbf{h}}_{r}\|\mathbf{W_{e}}\vec{\mathbf{e}}_{ir}]\big{)}\Big{)}}},$
where $\alpha_{ij}$ and $\vec{\mathbf{e}}_{ij}$ represent the attention weight and the edge features between $v_{i}$ and $v_{j}$ respectively, $\mathcal{N}_{i}$ denotes the neighbors of node $v_{i}$ in the graph $\mathcal{H}$, $\vec{\mathbf{h}}_{i}$ and $\vec{\mathbf{h}}_{j}$ are the representation vectors of node $v_{i}$ and $v_{j}$, and $\vec{\mathbf{a}}^{T}$, $\mathbf{W}$ and $\mathbf{W_{e}}$ are the learnable parameters. Compared to the attention mechanisms solely based on nodes (Veličković et al., 2017), the edge semantic attention mechanism fully considers the heterogeneity of edges by encoding various relationships into attention weights, enriching node representations. Furthermore, this mechanism empowers our model to dynamically assign weights to different edge types based on the specific communities, thereby preserving valuable information in diverse relationships.
In real-world heterogeneous information networks, not all networks have edge features. Therefore, we consider the following two scenarios:
  1. (1)
    The heterogeneous network includes edge features;
  2. (2)
    The heterogeneous network lacks edge features.
In the first scenario, where the heterogeneous information network encapsulates edge features, we can seamlessly apply equation 3. In the second scenario, to encode the information of edge types, we introduce an embedding matrix $\mathbf{E}$, which assigns a learnable feature vector to each edge type. Given =$\mathcal{E}=\{\phi(e_{i,j})|e_{i,j}\in E\}$ is the set of edge types in the graph $\mathcal{H}$, $\mathcal{E}_{t}$ denotes the $t$-th element in $\mathcal{E}$ and $\mathcal{E}_{t}=\phi(e_{i,j})$, we formalize the embedding matrix as follows:
(4) $\vec{\mathbf{e}}_{i,j}=\mathbf{E}[t,:]$
where $t\in\{1,2,\ldots,|\mathcal{E}|\}$ is a subscript of $\mathcal{E}$ indicating the corresponding type of edge $e_{i,j}$, $\mathbf{E}\in\mathbb{R}^{|\mathcal{E}|\times d_{e}}$, $|\mathcal{E}|$ denotes the number of edge types, $d_{e}$ denotes the feature dimension of edges. Through this approach, even in the absence of edge features, we are able to capture and utilize the intrinsic heterogeneity of the edges. This not only ensures the richness of the aggregate information but also enhances the adaptability of our algorithm to diverse HINs, thereby improving the overall effectiveness of multi-type community search.
4.3.2. Heterogeneous Encoder
The heterogeneous encoder is dedicated to capturing query-independent information, including node features as well as structural information in $\mathcal{H}_{h}$. Compared to previous studies in HINs (Jian et al., 2020; Fang et al., 2020; Jiang et al., 2022a; Zhou et al., 2023), the heterogeneous encoder dynamically considers the contributions of different relations to communities, rather than treating them equally.
The heterogeneous encoder is based on the idea of graph neural networks (GNN) (Zhang et al., 2019; Lv et al., 2021), equipped with the edge semantic attention mechanism, and iteratively combines the neighbor information of each node as follows:
[…]
Through Equation 5, we can uniformly consider structural information and node features with any dimension, avoiding ignoring node features or only considering low-dimensional features as in previous studies (Jian et al., 2020; Fang et al., 2021; Zhou et al., 2023). Compared to the vanilla GNN-based community search methods (Gao et al., 2021; Jiang et al., 2022b; Li et al., 2023) in homogeneous graphs, our proposed heterogeneous encoder integrates edge types and dynamically considers the contributions of different relations to communities, which significantly enhances the understanding of complex relationships in HINs.
4.3.3. Query Encoder
The query encoder is dedicated to obtaining query-dependent information in $\mathcal{H}_{q}$, including the local structural information around the query node, due to the query-dependent nature of community search. Compared to previous studies (Jian et al., 2020; Jiang et al., 2022b; Zhou et al., 2023), our proposed query encoder aggregates query information based on different types of edges to promote the model’s understanding of complex relationships in HINs. Specifically, the query encoder consists of multiple GNN layers, each incorporating the edge semantic attention mechanism as follows:
[…]
where $\vec{\mathbf{h}}_{qi}^{l}$ denotes the query-dependent representation of node $v_{i}$ in the $l$-th layer, and $\vec{\mathbf{h}}_{qi}^{0}=x_{q}[v_{i}]$. Note that for brevity, the learnable parameters, such as $\alpha_{ijk}^{l}$, $\mathbf{W}_{k}^{l}$ and $\mathbf{W_{r}}^{l}$, are no longer distinguished in Equations 5 and 6, but they train these parameters separately. The difference between Equation 5 and Equation 6 lies in their focus on different views, dealing with query-independent and query-dependent information, respectively. Through query encoder, we can better capture the local structural information around query nodes to facilitate community search.
4.3.4. Feature Fusion
The above heterogeneous encoder and query encoder respectively extract information from different views ($\mathcal{H}_{h}$ and $\mathcal{H}_{q}$). As the importance of each view may vary for each node, we cannot simply treat information from the heterogeneous encoder and query encoder equally. Therefore, we introduce a feature fusion module based on the attention mechanism after each layer to take full advantage of complementary information from different views. Specifically, node $v_{i}$ at layer $l$ has two different representations ($\vec{\mathbf{h}}_{i}^{l}$ and $\vec{\mathbf{h}}_{qi}^{l}$), originating from the heterogeneous encoder and the query encoder, respectively. Then, we use an attention mechanism to fuse these complementary representations for node $v_{i}$, as follows:
(7) $e_{i}^{l}=\vec{\mathbf{u}}^{\intercal}\cdot\vec{\mathbf{h}}_{i}^{l},$ $\quad e_{qi}^{l}=\vec{\mathbf{u}}_{q}^{\intercal}\cdot\vec{\mathbf{h}}_{qi}^{l},$
$\beta_{i}^{l}=\frac{\exp\left(e_{i}^{l}\right)}{\exp\left(e_{i}^{l}\right)+\exp\left(e_{qi}^{l}\right)},$ $\quad\beta_{qi}^{l}=\frac{\exp\left(e_{qi}^{l}\right)}{\exp\left(e_{i}^{l}\right)+\exp\left(e_{qi}^{l}\right)},$
$\vec{\mathbf{h}}_{fi}^{l}=\beta_{i}^{l}\cdot$ $\vec{\mathbf{h}}_{i}^{l}+\beta_{qi}^{l}\cdot\vec{\mathbf{h}}_{qi}^{l},$
where $\vec{\mathbf{h}}_{fi}^{l}$ is the fused node representation after $l$ layer, and $\vec{\mathbf{u}}^{\intercal}$ and $\vec{\mathbf{u}}_{q}^{\intercal}$ are the parameterized attention vectors, respectively. $\beta_{i}^{l}$ and $\beta_{qi}^{l}$ can be interpreted as the importance of different views for the node. As shown in Figure 2, for the query encoder, we use the fused node representation as the input for the next layer. For the heterogeneous encoder, we only use the output of the heterogeneous encoder itself in the $l$-th layer as the input of the $(l+1)$-th layer. In this way, we make the heterogeneous encoder independent of the query information, which provides stable prior knowledge about the heterogeneous information networks.
Furthermore, we use the output of the feature fusion module in the final layer as the node representation, which is utilized to compute the probability of each node belonging to a community, formalized as:
(8) $p_{i}=\sigma\Big{(}\text{MLP}(\vec{\mathbf{h}}_{fi}^{L})\Big{)},$
where $p_{i}$ denotes the probability of node $v_{i}$ belonging to the community, $\sigma$ indicates the Sigmoid function, and MLP refers to the multi-layer perceptron.
4.4. Search Algorithm
The search algorithm is designed to identify the target community based on user-specified node types $\mathcal{S}_{\mathcal{A}}$ and the probability of each node belonging to the community calculated by Equation 8. Moreover, in order to ensure the connectivity of communities, we design a search algorithm for multi-type communities in FCS-HGNN based on breadth-first search (BFS).
Input: User-specified node types $\mathcal{S}_{\mathcal{A}}$ in the target community, query node $q$, node probability $p_{i}$, $\mathcal{H}=(V,E,\psi,\phi)$, threshold $\gamma$
Output: The target community $\mathcal{C}_{q}$
1Initialize $\mathcal{Q}\leftarrow$ empty queue, $\mathcal{C}_{q}\leftarrow\{q\}$;
$\mathcal{V}\leftarrow\varnothing$ // recorded visited nodes
$\mathcal{Q}\leftarrow q$ // Add $q$ into the queue
2 while $\mathcal{Q}$ is not empty do
3       node $v_{i}\leftarrow$ dequeue from $\mathcal{Q}$;
4       if $v_{i}\notin\mathcal{V}$ then
             $\mathcal{V}\leftarrow\mathcal{V}\cup\{v_{i}\}$ // add visited nodes
5             foreach $v_{j}\in\mathcal{N}_{i}$ do
                   $\mathcal{Q}\leftarrow v_{j}$ // add new node for BFS
6                   if $\psi(v_{j})\in\mathcal{S}_{\mathcal{A}}$ and $p_{j}>\gamma$ then
                         // If the community conditions are satisfied
7                         $\mathcal{C}_{q}\leftarrow\mathcal{C}_{q}\cup\{v_{j}\}$;
8
9
10
11
return The target community $\mathcal{C}_{q}$.
Algorithm 1 Search algorithm for multi-type community
Algorithm 1 outlines our proposed search algorithm for the multi-type community. We first initialize a queue $\mathcal{Q}$ for BFS and a set $\mathcal{V}$ to record visited nodes (Lines 1-3). In the loop, for each neighbor $v_{j}$ of $v_{i}$, we do not restrict the criteria for adding $v_{j}$ to the queue $\mathcal{Q}$, as the neighbors of low probability ($p_{j}<\gamma$) or non-target type nodes ($\psi(v_{j})\notin\mathcal{S}_{\mathcal{A}}$) may also be community members (Line 9). On the contrary, for the node that satisfies the community conditions, we add it to the target community $\mathcal{C}_{q}$ (Lines 10-11). With Algorithm 1, we are able to output the target community based on the probability of each node belonging to the community, while ensuring the connectivity within the community.
4.5. Training and Online Query
4.5.1. Training
For the training stage, following previous work (Gao et al., 2021; Jiang et al., 2022b), we define the community search problem as a binary node classification problem, where the goal is to train the model to determine whether each node belongs to the community under query $q$. We denote $y\in\{0,1\}$ as the true label of each node under query $q$. Specifically, $y_{i}=1$ if and only if node $v_{i}$ belongs to community $\mathcal{C}_{q}$; otherwise, $y_{i}=0$. Then, we utilize the Binary Cross Entropy (BCE) function as the loss function to minimize the discrepancy between the predicted and ground-truth labels:
(9) $\mathcal{L}=-\sum_{q\in\mathcal{D}}\frac{1}{n}\sum_{i=1}^{n}\Big{(}y_{i}\log(p_{i})+(1-y_{i})\log(1-p_{i})\Big{)},$
where $y_{i}$ and $p_{i}$ are the ground-truth label and predicted probability of node $v_{i}$ under query $q$, respectively. The label of $y_{i}$ will change with different queries $q$. With the training data $\mathcal{D}$, we update the parameters of FCS-HGNN according to the Equation 9.
Input: Training data $\mathcal{D}$, HIN $\mathcal{H}$, the number of multi-head attention $K$, the number of layer $L$
Output: The optimal parameters of FCS-HGNN
1Initialize the parameters of FCS-HGNN;
2 while not converge do
3       for $(q,pos(q),neg(q))\in\mathcal{D}$ do
             /* Heterogeneous & Query Information Preprocessing */
4             $\mathcal{H}_{h}\leftarrow$ type-shared feature projection by Eq. 1;
5             $\mathcal{H}_{q}\leftarrow$ query-dependent information by Eq. 2;
             /* Information Integration */
6             $\vec{\mathbf{h}}^{1}\leftarrow\text{{Hetero. Encoder}}^{1}(\mathcal{H}_{h},\vec{\mathbf{h}}^{0})$ in Eq. 5;
7             $\vec{\mathbf{h}}_{q}^{1}\leftarrow\text{{Query Encoder}}^{1}(\mathcal{H}_{q},\vec{\mathbf{h}}_{q}^{0})$ in Eq. 6;
8             $\vec{\mathbf{h}}_{f}^{1}\leftarrow\text{{Feature Fusion}}^{1}(\vec{\mathbf{h}}^{1},\vec{\mathbf{h}}_{q}^{1})$ in Eq. 7;
9             $l=1$;
10             while $l<L$ do
11                   $\vec{\mathbf{h}}^{l+1}\leftarrow\text{{Hetero. Encoder}}^{l+1}(\mathcal{H},\vec{\mathbf{h}}^{l})$ in Eq. 5;
12                   $\vec{\mathbf{h}}_{q}^{l+1}\leftarrow\text{{Query Encoder}}^{l+1}(\mathcal{H}_{q},\vec{\mathbf{h}}_{f}^{l})$ in Eq. 6;
13                   $\vec{\mathbf{h}}_{f}^{l+1}\leftarrow\text{{Feature Fusion}}^{l+1}(\vec{\mathbf{h}}^{l+1},\vec{\mathbf{h}}_{q}^{l+1})$ in Eq. 7;
14                   $l=l+1$;
15
16            $p\leftarrow$ calculated probability of each node by Eq. 8;
17             Update the parameters according to the loss in Eq. 9;
18
19
return The optimal parameters of FCS-HGNN.
Algorithm 2 The training process of FCS-HGNN
Algorithm 2 outlines the training process of our proposed FCS-HGNN. In the loop, we first deal with the heterogeneity of nodes through type-shared feature projection to obtain $\mathcal{H}_{h}$ and construct the query-dependent graph $\mathcal{H}_{q}$ based on the query $q$ (lines 4-5). Then, we extract complementary information from different views through the heterogeneous encoder and query encoder, and integrate them via the feature fusion module (lines 6-14). Finally, we compute the probability of each node belonging to the community and update the model parameters (Lines 15-16). It is noteworthy that $\mathcal{H}_{h}$ and $\mathcal{H}_{q}$ differ solely in terms of node features, sharing the same graph structure. This characteristic ensures that there is no additional computational cost overhead incurred.
Furthermore, in our proposed FCS-HGNN, the single-type community is considered as a special case of the multi-type community, where $|\mathcal{S}_{\mathcal{A}}|=1$. During the training phase, the node types of the target community are agnostic to the model. This enables the model to focus on learning the inherent features of the HIN, query-dependent information, and the community patterns present in the training data. Then, we leverage the predicted probability of each node belonging to the community and the search algorithm to focus on the user-specified node types $\mathcal{S}_{\mathcal{A}}$. In other words, the training phase focuses on learning community patterns and the search algorithm considers user-specified node types. In this way, we are able to flexibly handle both single-type and multi-type communities, as shown in Table 1. Furthermore, our proposed FCS-HGNN significantly alleviates the burden of predefined metapaths and designing a series of relationship constraints.
4.5.2. Online Query
For the online query phase, we utilize the best model obtained during the training phase to execute queries instead of retraining the model for each query. For a new query $q_{new}$, we first employ the model to calculate the probability of each node belonging to the community. Then, we output the target community based on the search algorithm, as shown in Algorithm 1.
5. LARGE-SCALE FCS-HGNN
When dealing with large-scale graphs, traditional community search methods (Jiang et al., 2022a; Liu et al., 2023) usually improve efficiency by designing appropriate indexes. However, machine learning (ML)-based community search methods (Gao et al., 2021; Jiang et al., 2022b) face significant challenges when handling large-scale graphs due to the substantial computational costs. To address this issue, we introduce the neighbor sampling algorithm to enhance training efficiency and propose the depth-based heuristic strategy to improve query efficiency, resulting in the optimized algorithm named LS-FCS-HGNN.
5.1. Neighbor Sampling
To improve the training efficiency, we introduce a neighbor sampling algorithm, which recursively samples a fixed number of neighbors for each node to form a subgraph for training (Hamilton et al., 2017). Algorithm 3 outlines the neighbor sampling algorithm, where $S_{l}$ represents the set of nodes sampled in the $l$-th layer, $N_{v_{i}}^{l-1}$ represents the set of nodes sampled from the neighbors of $v_{i}$, utilized for training in the $(l-1)$-th layer, and fanouts $\{f_{1},f_{2},...,f_{L}\}$ are hyperparameters indicating the number of neighbors that should be sampled at each layer. We initialize $S_{L}$ with the labeled nodes, as the probabilities of these nodes are ultimately used to compute the loss in Equation 9 (Line 1). Then, for each node $v_{i}$ in $S_{l}$, we sample a fixed number $f_{l-1}$ of nodes from the neighbors of $v_{i}$, and add them to $S_{l-1}$ (Lines 2-6). Finally, we extract the corresponding subgraph $\mathcal{H}^{\prime}$ from $\mathcal{H}$ based on the nodes sampled at each layer (Line 7).
Input: labeled Nodes ($q$, $pos(q)$ and $neg(q)$), $\mathcal{H}$, the number of layers $L$, Fanouts $f_{1},f_{2},...,f_{L}$
Output: The subgraph for training
$S_{L}\leftarrow\{q,pos(q),neg(q)\}$// Initialization
1 for $l=L$ to $1$ do
       $S_{l-1}\leftarrow\varnothing$ // $S_{l-1}$ record the nodes sampled in the $(l-1)$-th layer
2       foreach node $v_{i}$ in $S_{l}$ do
3             $N_{v_{i}}^{l-1}\leftarrow$ Sample $f_{l-1}$ neighbors of $v_{i}$ from $\mathcal{H}$;
4             $S_{l-1}\leftarrow S_{l-1}\cup\{N_{v_{i}}^{l-1}\}$;
5
6
7Extract the subgraph $\mathcal{H}^{\prime}$ of $\mathcal{H}$ with $\bigcup_{l=0}^{L}S_{l}$;
return The subgraph $\mathcal{H}^{\prime}$
Algorithm 3 Neighbor Sampling of LS-FCS-HGNN
Input: Training data $\mathcal{D}$, HIN $\mathcal{H}$, the number of multi-head attention $K$, the number of layer $L$, Fanouts $f_{1},f_{2},...,f_{L}$
Output: The optimal parameters of LS-FCS-HGNN
1Initialize the parameters of LS-FCS-HGNN;
2 while not converge do
3       for $(q,pos(q),neg(q))\in\mathcal{D}$ do
             /* Neighbor Sampling */
4             $\mathcal{H^{\prime}}\leftarrow$ Algorithm 3 ($\{q,pos(q),neg(q)\}$, $\mathcal{H}$, $L$, $f_{1},f_{2},...,f_{L}$);
             /* Training model */
5             Line 4-16 in Algorithm 2 with $\mathcal{H^{\prime}}$;
6
7
return The optimal parameters for LS-FCS-HGNN.
Algorithm 4 The training process of LS-FCS-HGNN
In addition, based on the neighbor sampling (Algorithm 3), we outline the training process of LS-FCS-HGNN in Algorithm 4. Compared to Algorithm 2, LS-FCS-HGNN incorporates the process of neighbor sampling (Line 4) and utilizes $\mathcal{H}^{\prime}$ instead of $\mathcal{H}$ for subsequent training (Line 5). It is worth noting that we perform neighbor sampling at each epoch to obtain a variety of subgraphs, which allows the model to observe various structures and thereby comprehensively understand the entire community pattern. Therefore, the neighbor sampling strategy effectively reduces the computational costs of community search in large-scale graphs while maintaining the community effectiveness. Subsequent experiments in Section 6.2.3 further demonstrate the effectiveness and efficiency of the neighbor sampling strategy.
5.2. Depth-based Heuristic Strategy
To improve the query efficiency in the large-scale graphs, we propose the depth-based heuristic search strategy. The core hypothesis of our proposed depth-based heuristic strategy is that: the nodes closer to the query node are more likely to belong to the same community. This assumption is grounded in real-life observations. In many cases, individuals within a community tend to have stronger connections with each other compared to individuals outside the community. In other words, nodes within the same community typically have more edges and shorter paths between them. Therefore, we propose the depth-based heuristic strategy to limit the exploration depth, reducing the search on distant nodes and thereby enhancing search efficiency.
Algorithm 5 outlines the depth-based heuristic search algorithm. Compared to Algorithm 1, the queue $\mathcal{Q}$ not only stores node information but also records the depth of each node from the query node $q$ (Line 3). Since the depth explored by BFS is always the closest distance from the node to the query node, we use the depth to determine whether to explore the node (Line 6). Then, we do not restrict the enqueue conditions and add a node to the target community if the node satisfies the community conditions (Lines 9-11).
Input: User-specified node types $\mathcal{S}_{\mathcal{A}}$ in the target community, query node $q$, node probability $p_{i}$, $\mathcal{H}=(V,E,\psi,\phi)$, threshold $\gamma$, maximum depth $d_{max}$
Output: The target community $\mathcal{C}_{q}$
1Initialize $\mathcal{Q}\leftarrow$ empty queue, $\mathcal{C}_{q}\leftarrow\{q\}$;
$\mathcal{V}\leftarrow\varnothing$ // recorded visited nodes
$\mathcal{Q}\leftarrow(q,0)$ // Add $q$ and depth 0 into the queue
2 while $\mathcal{Q}$ is not empty do
3       (node $v_{i}$, depth $d$) $\leftarrow$ dequeue from $\mathcal{Q}$;
4       if $v_{i}\notin\mathcal{V}$ and $d<d_{max}$ then
             $\mathcal{V}\leftarrow\mathcal{V}\cup\{v_{i}\}$ // add visited nodes
5             foreach $v_{j}\in\mathcal{N}_{i}$ do
                   $\mathcal{Q}\leftarrow(v_{j},d+1)$ // add new node for BFS
6                   if $\psi(v_{j})\in\mathcal{S}_{\mathcal{A}}$ and $p_{j}>\gamma$ then
7                         $\mathcal{C}_{q}\leftarrow\mathcal{C}_{q}\cup\{v_{j}\}$;
8
9
10
11
return The target community $\mathcal{C}_{q}$.
Algorithm 5 Depth-based heuristic search algorithm
5.3. Time Complexity
In the following, we analyze the time complexity of our proposed algorithms to analyse the efficiency of FCS-HGNN and LS-FCS-HGNN.
Theorem 1 ().
Let $T$ denote the number of training iterations, $|V|$ represent the number of nodes, $|E|$ represent the number of edges, $|\mathcal{D}|$ be the quantity of the training set, and $F$ and $F^{\prime}$ represent the feature dimensions before and after linear transformation, respectively. The time complexity for a layer of heterogeneous encoder or query encoder is $\mathcal{O}(|V|FF^{\prime}+|E|F^{\prime})$, where $|V|FF^{\prime}$ is the time complexity for the transformation of feature vectors for all nodes, and $|E|F^{\prime}$ is the time complexity for computing attention. For the FCS-HGNN with $L$ layers, the time complexity for each iteration would be $\mathcal{O}(L(|V|FF^{\prime}+|E|F^{\prime}))$. Thus, the time complexity of the training process of FCS-HGNN (Algorithm 2) is $\mathcal{O}(TL|\mathcal{D}|(|V|FF^{\prime}+|E|F^{\prime}))$, which primarily depends on the size of the graph, such as $|V|$ and $|E|$.
Theorem 2 ().
Let $|V|$ and $|E|$ represent the number of nodes and edges respectively, the time complexity of search algorithm (Algorithm 1) is $\mathcal{O}(|V|+|E|)$. Since we only need to perform a forward propagation during the inference phase, the time complexity of the query process for FCS-HGNN is $\mathcal{O}(L(|V|FF^{\prime}+|E|F^{\prime})+|V|+|E|)$.
Theorem 3 ().
Let LS-FCS-HGNN have $L$ layers and there are $n_{l}$ nodes to be sampled in the $l$-th layer. The time complexity of neighbor sampling (Algorithm 3) is $\mathcal{O}(L\sum_{l=1}^{L}n_{l})$, which primarily depends on the number of operations ($n_{l}$) performed in each layer.
Theorem 4 ().
Let $n$ and $e$ represent the total number of nodes and edges in the subgraph respectively, where $n=\sum_{l=1}^{L}n_{l}$. According to the theorem 1 and 3, the time complexity of the training process of LS-FCS-HGNN (Algorithm 4) is $\mathcal{O}(TL|\mathcal{D}|(nFF^{\prime}+eF^{\prime}+Ln))$, where $Ln$ is the time complexity of neighbor sampling.
Theorem 5 ().
The time complexity of depth-based heuristic search algorithm (Algorithm 5) is bounded by $\mathcal{O}(|V|+|E|)$, which depends on the maximum exploration depth $d_{max}$ as well as the graph structure. Thus, the time complexity of the query process of LS-FCS-HGNN is bounded by $\mathcal{O}(L(nFF^{\prime}+eF^{\prime}+Ln)+|V|+|E|)$.
Since $n$ and $e$ are much smaller than $|V|$ and $|E|$, the training and query efficiency of LS-FCS-HGNN will be significantly higher than that of FCS-HGNN. In the next section, we compare the time performance of our proposed algorithm with state-of-the-art methods.
6. EXPERIMENTS
In this section, we first elaborate on the experimental setup. Subsequently, we compare the effectiveness and the efficiency of our proposed methods against state-of-the-art methods in both single-type and multi-type community searches. Finally, we conduct extensive ablation studies and parameter sensitivity analyses on our proposed methods.
6.1. Experimental Setup
6.1.1. Datasets.
To comprehensively evaluate the performance of our proposed method, we conducted experiments on five real-world heterogeneous information networks, i.e., IMDB (Luo et al., 2021), DBLP (Fu et al., 2020), ACM (Luo et al., 2021), Freebase (Lv et al., 2021), and OGB-MAG (Hu et al., 2020). Table 2 reports the statistics of the datasets, where $|\mathcal{A}|$ and $|\mathcal{E}|$ denote the number of node types and edge types, respectively. Node features are necessary for all datasets. Notably, Freebase possesses a wealth of edge relationships, while OGB-MAG is a large-scale HIN derived from the OGB benchmark (Hu et al., 2020).
Table 2. Dataset statistics
Dataset $|V|$ $|E|$ $|\mathcal{A}|$ $|\mathcal{E}|$
IMDB 19,103 74,880 4 6
DBLP 26,128 265,694 4 6
ACM 31,807 160,321 4 6
Freebase 180,098 1,057,688 8 36
OGB-MAG 1,939,743 36,805,743 4 7
6.1.2. Community Generation.
In reality, it is challenging to manually collect labeled communities for training, especially multi-type communities in large-scale graphs. Thus, we adopt traditional community search methods in HINs to automatically label the single-type (Jiang et al., 2022a) or multi-type communities (Jian et al., 2020) as the ground truth. Specifically, the objective of (Jiang et al., 2022a) is to identify $(k,\Psi)$-NMC communities containing nodes of a single type, while (Jian et al., 2020) aims to search for relational communities containing nodes of different types. Since other traditional community search methods (Fang et al., 2020; Zhou et al., 2023) in HINs focus on different community patterns that are not consistent with those of (Jian et al., 2020) and (Jiang et al., 2022a), it would be unfair to compare these methods. However, machine learning (ML)-based methods do not rely on predefined community patterns; instead, they learn community patterns in a data-driven manner. Therefore, although these communities (Jian et al., 2020; Jiang et al., 2022a) impose strict constraints on community patterns, they provide a relatively fair comparison for ML-based methods in the absence of labeled communities in HINs. For these reasons, in this paper, we mainly compare with state-of-the-art ML-based methods. We randomly select one node from the groundtruth community as a query node $q$, randomly select three nodes as $pos(q)$, and three nodes outside of the groundtruth community as $neg(q)$. In order to evaluate the effectiveness of the algorithm for unseen queries, we partition the training, validation, and test sets with a ratio of 1:1:3, which minimizes the labeling information. For the single-type community, we designate one type of node as the primary node. For the DBLP dataset, we consider either author or paper nodes as the primary nodes, resulting in two variants: DBLP-A and DBLP-P, respectively. For the other datasets, we designate the primary node for single-type community according to (Luo et al., 2021). For the multi-type community, we set the paper and author nodes as primary nodes for DBLP, ACM, and OGB-MAG, movie and actor nodes as primary nodes for IMDB, film and people nodes as primary nodes for Freebase.
6.1.3. Baselines.
To our knowledge, we are the first to propose the multi-type community search method based on graph neural networks in HINs, thus there are no direct studies available for comparison. Therefore, we compare our proposed FCS-HGNN and LS-FCS-HGNN with state-of-the-art methods in homogeneous graphs, including ICS-GNN (Gao et al., 2021), QD-GNN (Jiang et al., 2022b), COCLE and COCLEP (Li et al., 2023). Specifically, we transform HINs into homogeneous graphs using meta-paths (Wang et al., 2019; Zhang et al., 2019; Luo et al., 2021) and augment these baselines with our proposed type-shared feature projection to enable these methods to identify multi-type communities. Other methods (Hashemi et al., 2023; Behrouz and Hashemi, 2022; Fang et al., 2023) may face challenges in being applied to HINs even with preprocessing, therefore we do not consider comparing with them in this paper.
6.1.4. Evaluation Metrics.
To evaluate the quality of the identified communities, we employed three metrics: F1-score (Fang et al., 2023), Jaccard Similarity (Zhang et al., 2020), and Normalized Mutual Information (NMI) (Danon et al., 2005). These metrics are designed to assess different aspects of the identified communities (Li et al., 2023), aiding in comprehensive evaluation of various methods.
6.1.5. Implementation Details.
For our model, we construct $L=2$ layers with dimensions of 64 or 128, and used Adam (Kingma and Ba, 2014) for optimization. As for other hyperparameters, we set the learning rate to 0.001, the regularization parameter to 1e-4, the dropout rate to 0.5, the number of attention heads $K$ to 8, and the fanouts to $f_{1}=20,f_{2}=10$. The threshold $\gamma$ is obtained by searching for the best F1 score from the validation set. For LS-FCS-HGNN, we traverse the maximum depth $d_{max}$ from 2 to 10 and record the best results. For all the baselines, we fine-tuned their parameters according to their original papers and recorded their best performance. All experiments were conducted on Ubuntu 20.04 equipped with 2.40 GHz Intel(R) Xeon(R) Gold 6240R CPUs, 512GB memory, and 4 NVIDIA Tesla V100 GPUs.
6.2. Community Search Performance
In this section, we conducted a comprehensive comparison with state-of-the-art methods for both single-type and multi-type community searches in HINs. Then, we compare the training and query efficiency with state-of-the-art methods and conduct a case study.
Table 3. Performance comparison in single-type community search (in percentage).
Method DBLP-A DBLP-P ACM IMDB Freebase OGB-MAG
F1 JS NMI F1 JS NMI F1 JS NMI F1 JS NMI F1 JS NMI F1 JS NMI
ICS-GNN 58.38 46.94 37.44 73.75 68.88 69.16 42.48 31.22 25.23 31.2 19.42 8.21 46.44 43.82 45.07 o.o.m o.o.m o.o.m
QD-GNN 54.43 42.84 27.34 66.81 65.42 64.75 40.43 26.33 21.39 61.01 44.46 10.47 47.48 39.2 37.58 o.o.m o.o.m o.o.m
COCLE 57.70 43.63 31.59 68.12 62.02 67.99 40.69 26.98 34.04 60.47 44.19 13.44 38.95 33.14 36.00 o.o.m o.o.m o.o.m
COCLEP 63.41 48.64 35.05 66.69 59.23 60.89 55.76 41.74 38.35 50.93 34.53 10.53 43.8 38.19 36.62 17.85 11.33 9.71
FCS-HGNN 77.45 64.38 62.28 78.99 74.23 74.59 76.73 75.38 75.52 67.35 51.07 42.29 54.44 51.82 53.48 o.o.m o.o.m o.o.m
LS-FCS-HGNN 75.69 63.65 62.27 81.30 76.35 76.56 79.39 76.32 75.81 68.70 52.61 42.48 54.44 51.82 53.48 24.70 15.90 11.75
Table 4. Performance comparison in multi-type community search.
Method DBLP ACM IMDB Freebase OGB-MAG
F1 JS NMI F1 JS NMI F1 JS NMI F1 JS NMI F1 JS NMI
ICS-GNN 40.40 27.22 35.02 37.48 30.68 34.14 35.82 28.41 22.04 31.47 22.00 26.86 o.o.m o.o.m o.o.m
QD-GNN 46.78 39.57 42.74 38.23 33.69 35.28 39.55 33.93 28.45 34.96 27.10 26.38 o.o.m o.o.m o.o.m
COCLE 52.82 41.67 45.23 41.99 28.89 33.19 41.39 32.68 35.31 33.16 21.70 22.15 o.o.m o.o.m o.o.m
COCLEP 51.30 40.87 45.14 37.13 27.49 31.67 36.85 29.27 32.88 56.71 46.27 47.64 11.73 9.42 5.77
FCS-HGNN 55.45 42.13 47.48 53.01 40.88 46.31 44.11 36.80 40.64 68.07 54.50 57.78 o.o.m o.o.m o.o.m
LS-FCS-HGNN 55.19 41.87 47.15 52.67 40.43 45.86 44.15 36.82 40.69 89.00 84.54 85.39 19.07 11.85 8.45
6.2.1. Single-type Community Search.
Since the single-type community is a special case of the multi-type community, we first compare the effectiveness of our proposed algorithms in the context of the single-type community search. Table 3 displays the experimental results of single-type community search in HINs. The best results from each dataset are highlighted in bold, while the best results from the baseline method are underlined for comparison. We also indicate the case of ”out of memory” (o.o.m).
From Table 3, it is evident that our proposed methods outperform other state-of-the-art methods across all datasets, achieving average improvements of 11.08%/13.08%/18.68% on the F1/JS/NMI metrics, respectively. This demonstrates the effectiveness of our methods when identifying single-type community search in HINs. We observe that these baselines generally perform poorly on the NMI metric, indicating that they fail to capture the structure and node distribution of the target communities adequately. However, our method not only significantly surpasses the baselines on the NMI metric but also does not exhibit lopsided performance across the three metrics. This indicates that our proposed methods effectively learn the complex community structure in HINs, which is crucial for community search. Specifically, our methods can extract valuable information from two different views $\mathcal{H}_{h}$ and $\mathcal{H}_{q}$ and dynamically integrate information through the feature fusion module. In each encoder, our proposed edge semantic attention mechanism effectively helps the model learn the complex community structure in the HIN during the information integration. Furthermore, we perform a comparative analysis between FCS-HGNN and LS-FCS-HGNN. It is illuminating to observe that LS-FCS-HGNN produced superior results in most datasets except for DBLP-A, demonstrating the effectiveness of the neighbor sampling and depth-based heuristic strategies. Under the guidance of the depth-based heuristic strategy, LS-FCS-HGNN eliminates the interference from distant nodes, achieving superior results compared to FCS-HGNN which employs a full-graph search. Simultaneously, this validates the reasonableness of our hypothesis that nodes closer to the query node are more likely to belong to the same community.
6.2.2. Multi-type Community Search.
Then, we evaluate the effectiveness of our proposed methods in the context of multi-type community search. Table 4 presents the experimental results of multi-type community search in HINs. It is worth emphasizing that for our proposed methods, the transition from single-type community search to multi-type community search only requires modifying the user-specific node types $\mathcal{S}_{\mathcal{A}}$ of their desired communities, which is very easy for users. Conversely, we undertake a considerable amount of work to adapt these baselines for multi-type community search. There is no need for any changes to the graphs or models for our methods, which significantly improves the flexibility and applicability of the algorithm.
As shown in Table 4, our proposed methods exhibit outstanding performance across all datasets, achieving average improvements of 11.21%/10.24%/11.81% on the F1/JS/NMI metrics, respectively. On one hand, the type-shared feature projection adeptly addresses the heterogeneity of nodes in the absence of meta-paths or relational constraints, which maps nodes from different feature spaces into a unified feature space, thereby enhancing information integration. On the other hand, the edge semantic attention mechanism selectively aggregates information from neighbors based on different relations, effectively leveraging fine-grained heterogeneous information and facilitating the model’s understanding of HINs. Our subsequent ablation experiments further confirmed this observation. Moreover, we are able to obtain similar conclusions to single-type community search: our proposed neighbor sampling (Algorithm 3) and depth-based heuristic search strategies (Algorithm 5) not only empower LS-FCS-HGNN to efficiently identify communities in large-scale graphs, but also improve the quality and precision of community identification. This further demonstrates the effectiveness of our methods.
6.2.3. Training Efficiency.
Figure 3 illustrates the training time per epoch for various methods on all datasets. As ICS-GNN requires retraining for each query, we did not record its result of training efficiency. On the OGB-MAG dataset, QD-GNN, COCLE, and FCS-HGNN run out of GPU memory. As shown in Figure 3, we observe that when the graph size is small, the training efficiency of each method is similar, but as the size increases, the differences in training efficiency become more apparent. In general, FCS-HGNN exhibits better training efficiency than COCLE and COCLEP. However, our proposed LS-FCS-HGNN demonstrates exceptional training efficiency, surpassing all competing methods. Especially as the graph size increases, the advantage of LS-FCS-HGNN’s training efficiency becomes increasingly significant.
6.2.4. Query Efficiency.
In addition to training efficiency, we also thoroughly compared the query efficiency of each method, as shown in Figure 4. It is worth noting that each grid on the y-axis of Figure 4 differs by a factor of 10. Since ICS-GNN retrains the whole model for each new query, it exhibits the lowest query efficiency. Similar to training efficiency, the query efficiency of FCS-HGNN significantly outperforms that of COCLE and COCLEP. However, with the assistance of our proposed depth-based heuristic search strategy, the query efficiency of LS-FCS-HGNN is significantly superior to all other methods. As the graph size continues to increase, this advantage becomes even more pronounced. This further demonstrates the effectiveness of our proposed depth-based heuristic search strategy: reducing exploration of distant nodes to enhance query efficiency while maintaining or even improving query accuracy (as shown in Table 3 and 4).
Figure 3. Training Efficiency.
Figure 4. Query Efficiency.
6.2.5. Case Study
We conducted a case study on the DBLP dataset to compare with traditional community search methods (Jiang et al., 2022a; Jian et al., 2020) and COCLEP (Li et al., 2023). In Figure 5, figures (a-c) depict cases of single-type communities, while (d-e) represent cases of multi-type communities. specifically, Figure 4(a) is a case of ground truth in single-type community by (Jian et al., 2020) with $k=3$ and $\Psi=\{(APTPA)\}$. Figure 4(d) is a case of ground truth in multi-type community by (Jiang et al., 2022a) with relational constraint $\{\text{author},\text{paper},k=3\}$. In Figure 5, the query node is marked in white color, the black nodes are the nodes in ground-truth community, and the red nodes represent nodes returned by other methods but not present in the ground-truth communities. In nodes of different shapes, circles represent authors, while squares represent papers. We observe that our method returns a community that is the closest to the ground-truth community. Compared to COCLEP, our method demonstrates a better understanding of the complex relationships and community information in HINs.
Refer to caption (a) groundtruth (Jiang et al., 2022a) Refer to caption (b) COCLEP Refer to caption (c) LS-FCS-HGNN Refer to caption (d) groundtruth (Jian et al., 2020) Refer to caption (e) COCLEP Refer to caption (f) LS-FCS-HGNN
Figure 5. Case Study.
6.3. Ablation Study and Parameter Sensitivity Analysis
In this section, we conduct comprehensive ablation studies to assess the effectiveness of each module, and perform the sensitivity analyses on the key hyperparameters.
Figure 6. Ablation Study of EA.
Figure 7. Ablation Study of Query Enc.
Refer to caption (a) DBLP Refer to caption (b) ACM
Refer to caption (a) DBLP Refer to caption (b) ACM
Refer to caption (a) DBLP Refer to caption (b) ACM
Figure 6. Ablation Study of EA.
Figure 7. Ablation Study of Query Enc.
Figure 8. Ablation Study of Hetero. Enc.
Figure 9. Parameter Analysis of $d_{max}$.
Figure 10. Parameter Analysis of $f_{0}$.
Refer to caption (a) DBLP Refer to caption (b) ACM
Refer to caption (a) DBLP Refer to caption (b) ACM
Refer to caption (a) DBLP Refer to caption (b) ACM
Figure 9. Parameter Analysis of $d_{max}$.
Figure 10. Parameter Analysis of $f_{0}$.
Figure 11. Parameter Analysis of $f_{1}$.
6.3.1. Ablation Study of Edge Semantic Attention Mechanism.
In our proposed methods, the edge semantic attention mechanism is a crucial component of both the heterogeneous encoder and query encoder. Therefore, we conduct ablation studies to evaluate the effectiveness of the edge semantic attention mechanism in community search. We substitute the edge semantic attention with the attention mechanism mentioned in GAT (Veličković et al., 2017) which only depends on nodes, and the results are shown in Figure 8, where ’wo’ is an abbreviation for without, and ’EA’ is an abbreviation for edge semantic attention mechanism. It is evident that removing the edge semantic attention leads to a significant decrease in performance, even while retaining the node attention mechanism. This suggests that edge heterogeneity is crucial for community search tasks in HINs. Our method effectively captures and utilizes the heterogeneous information present in different edges of HINs, thereby enhancing the model’s understanding of the graph structure.
6.3.2. Ablation Study of Query and Heterogeneous Encoders.
We investigate the query and heterogeneous encoder in FCS-HGNN to gain a deeper understanding of their individual contributions to the overall performance. As illustrated in Figure 8, the impact of the query encoder on the accuracy of community search is relatively small, as the formation of communities primarily relies on the inherent features in HINs. However, the incorporation of the query encoder proves valuable by eliminating the need for retraining, thereby enhancing the efficiency of online queries. As illustrated in Figure 8, the heterogeneous encoder is crucial for community search, and its absence brings significant performance degradation. Without the heterogeneous encoder, the model relies solely on graph structure and query information to predict communities. Since our ground truth communities are labeled by traditional community search methods (Jiang et al., 2022a; Jian et al., 2020), which typically rely only on graph structure, the experimental results in Figure 8 still maintain relatively high accuracy.
6.3.3. Parameter Analysis of Maximum Exploration Depth $d_{max}$.
We investigate the key parameter $d_{max}$ in Algorithm 5, which affects the depth of exploration. Figure 11 displays the performance with varying $d_{max}$. We observe that as $d_{max}$ increases, the metrics initially increase rapidly and then tend to stabilize. The optimal $d_{max}$ varies with the different datasets. Overall, due to the ”Six Degrees of Separation” theory (Guare, 2016), $d_{max}$ around 6 can achieve the best performance. Moreover, there is not a significant performance decline as $d_{max}$ increases. Therefore, the setting of $d_{max}$ is not challenging.
6.3.4. Parameter Analysis of fanouts $f_{0}$ and $f_{1}$.
Then, we investigate the impact of the hyperparameter fanouts $f_{0}$ and $f_{1}$ in Algorithm 3. For the analyses, $f_{0}$ is varied within {5,10,15,20} with a controlled $f_{1}=10$, and vice versa, where $f_{1}$ is varied within the same range while maintaining $f_{0}=20$. Figure 11 and 11 illustrate the performance with varying $f_{0}$ or $f_{1}$. We observe that the model is not sensitive to the fanouts $f_{0}$ or $f_{1}$, which is consistent with our previous analysis. Due to the random sampling of neighbors in each epoch, the model is able to learn node features and community structure in an efficient and stable manner.
7. CONCLUSIONS
In this paper, we propose a novel query-driven heterogeneous graph neural network, named FCS-HGNN, for multi-type community search in HINs. Compared to existing methods, it can flexibly identify both single-type and multi-type communities without relying on predefined meta-paths or user-specific relationship constraints, significantly alleviating the burden on users. FCS-HGNN dynamically consider the contribution of each relation instead of treating them equally, thus capturing more fine-grained heterogeneous information. Furthermore, to improve efficiency on large-scale graphs, we introduce a neighbor sampling algorithm to enhance training efficiency and propose the depth-based heuristic search strategy to improve query efficiency. We thoroughly evaluate our proposed methods on five real-world HINs and demonstrate its significant superiority in terms of effectiveness and efficiency.
Acknowledgements.
We would like to express our gratitude to the teams led by Lei Chen (Jian et al., 2020) and Yixiang Fang (Jiang et al., 2022a) for laying the groundwork with their outstanding research achievements in this field. Fangda Guo is the corresponding author.